OpenStudy (anonymous):

What is the slope of the line given by the equation y + 2 = −3(x − 7)?

OpenStudy (campbell_st):

well the slope is the value outside the brackets

OpenStudy (anonymous):

Well, this is the same equation from earlier so we have y = -3x +19 from there. The slope is the m value is the equation y = mx+b so the m value would be -3 therefore the slope is -3

OpenStudy (campbell_st):

the point slope formula is \[y - y_{1} = m(x - x_{1})\] m = slope